OpenStudy (anonymous):

A unit vector orthogonal to both a=<−2, 0, −1> and b=i−j−k is u= ?

OpenStudy (amistre64):

the cross of the 2

OpenStudy (anonymous):

so that gives me <-1,-2, 2>, but that's wrong?

OpenStudy (amistre64):

-2 0 -1 1 -1 -1 -------- 0(-1)-(-1)(-1) = -1 -1(1)- is your cross right? because thats just the vector to it; you have to then divide out the magnitude to get it to a unit size

OpenStudy (anonymous):

and then my mag. is 3, then?

OpenStudy (anonymous):

so all the values i listed in my cross product, divided by 3?

OpenStudy (amistre64):

lets assume your cross is right and go with it :) sqrt(1+4+4) = sqrt(9) = 3, yes

OpenStudy (amistre64):

divide out the 3 to get it to unit length, a length of 1

OpenStudy (amistre64):